# Low-Altitude Wireless Networks: A Survey

### v1

@@ 2.2 LAWN Channel Modeling
Channel modeling serves as the cornerstone for ensuring safe and reliable UAV operations, as it captures the propagation behaviors that underpin performance evaluation and system design. In this context, spectrum allocations define the operating conditions that directly shape the propagation characteristics. To this end, the International Civil Aviation Organization (ICAO) requires that control and non-payload communication (CNPC) links operate in protected spectrum bands. In line with this requirement, the International Telecommunication Union (ITU) has designated specific portions of the L-band (960–977 MHz) and C-band (5030–5091 MHz) for CNPC services in UAV systems [30]. For beyond-line-of-sight (LoS) operations involving satellite communication, the Ku-band and Ka-band have been reserved under the aeronautical safety spectrum framework to support robust satellite-based CNPC connectivity. In addition, national regulators, such as those in China, have authorized supplementary LoS frequency bands to meet the rising demand for UAV deployments in densely utilized low-altitude airspaces [31]. Based on the allocated bands, traditional studies on aeronautical channel modeling primarily target high-altitude, long-range aircraft operating under transcontinental or military flight conditions. The work in [32] introduced one of the earliest statistical frameworks for such environments, analyzing Doppler spread and multipath dispersion caused by rapid airborne motion. In [33], the authors categorized modeling approaches into deterministic, stochastic, and geometry-based types, highlighting the influence of aircraft altitude, elevation angle, and antenna orientation on channel characteristics such as path loss and Ricean K-factors. The authors in [34] further examined wireless datalink requirements for large-scale unmanned systems, analyzing the trade-offs among various candidate technologies under high-mobility and limited-spectrum conditions. While primarily focused on conventional high-altitude aircraft, the resulting models and statistical parameters offer a foundational reference for extending channel characterization into LAWNs.
@@
The A2G channel is pivotal for facilitating data exchange within LAWNs, making accurate modeling of the A2G propagation channel essential for the design of resilient drone-based communication systems. In urban environments where LAWNs are commonly deployed, the propagation channels are often shaped by dense building structures, irregular street canyons, and pronounced multi-path effects, leading to significant challenges for traditional propagation models. A representative statistical modeling approach is presented in [35], where the authors conducted extensive UAV field measurements at low elevation angles across urban environments. They developed a time-series generator capable of capturing the elevation-dependent fading characteristics, providing a realistic stochastic representation of mobile A2G links for simulation and system design. Furthermore, the authors in [36] studied 6G low-altitude mmWave massive multiple-input multiple-output (MIMO) A2G channels in urban scenarios, capturing space-time-frequency non-stationarity with UAV-specific parameters. To further quantify large-scale shadowing behavior of A2G channels, an elevation-dependent shadowing model was proposed in [37]. Originally developed for high-altitude platforms (HAPs), the model remains applicable to LAWNs operating over urban areas. The key insight is that the shadowing loss exhibits a strong dependence on the elevation angle, following an exponential trend as the UAV moves closer to the horizon, thus affecting LoS probability and average signal strength. The work of [38] investigated an analytical A2G path loss model that integrates stochastic geometry with ray-tracing results, which separates LoS and non-LoS (NLoS) components and incorporates a probabilistic LoS function based on elevation angle and building distribution. To handle fast-varying urban scenarios, a hybrid learning-based channel prediction method is proposed in [39], where the authors combined model-based signal evolution with data-driven learning to forecast near-future channel states for UAV relays navigating complex urban spaces.
@@
In addition to urban deployments, the study in [40] proposed an analytical path-loss model for suburban A2G transmissions. By incorporating UAV height and LoS probability, it facilitates efficient planning of UAV trajectories and resource allocation in semi-urban environments. Regarding over-water scenarios, the authors in [41] developed a geometry-based multi-path model capturing sea surface reflections and long-range propagation delays for A2G links. Meanwhile, the work in [42] conducted C-band measurements over sea surfaces at low altitudes, characterizing channel fading patterns and multi-path effects essential for UAV telemetry and maritime missions. Notably, LAWN systems are increasingly integrated into cellular networks, which exhibit distinct propagation characteristics. The work in [43] conducted empirical studies of UAV communication over commercial long-term evolution (LTE) networks, which revealed that drones flying above the downtilted beam of base station antennas are likely to result in higher path loss and interference levels. The authors in [44] further explored mid-band 5G service delivery for ultra-low-altitude UAV base stations, whose measurements demonstrated significant Doppler effects and elevation-dependent delay profiles, emphasizing the importance of precise modeling for 5G aerial deployments. To complement theoretical modeling, a series of wideband channel measurement campaigns were conducted in [45, 46, 47], targeting A2G links over diverse terrains including over-water, mountainous, and suburban areas, providing empirical models capturing delay spread, angular dispersion, and Ricean fading statistics with varying antenna heights and aircraft velocities.
@@ 3.1 Integrated Sensing and Communications
ISAC represents a transformative paradigm that unifies wireless communications and environmental sensing into a single framework [50, 51, 52, 53, 54, 55]. By adopting a co-design architecture, ISAC eliminates the need for deploying separate sensing or communication subsystems on aerial platforms within LAWNs. This integration not only reduces the spectral footprint, hardware redundancy, and energy consumption, but also enhances the agility of the system. For instance, a UAV transmitting communication signals to ground users can simultaneously leverage the backscattered echoes for obstacle detection or trajectory tracking of neighboring UAVs [56], thereby avoiding reliance on dedicated radar or LiDAR sensors.
@@
  • Stage 2 (Collaborative ISAC systems): In this stage, each UAV integrates sensing and communication functionalities into a unified platform by sharing hardware and spectral resources. This design improves spectral efficiency, reduces payload burden, and enhances cost-effectiveness compared to separated systems. However, its benefits remain limited to individual UAVs, and scalability across large networks is not yet fully achieved.
@@
This evolutionary trajectory highlights the maturation of ISAC from a dual-functional system into a network-wide intelligent infrastructure, which is particularly crucial for supporting the burgeoning LAE. In such scenarios, UAV networks serve as a vital enabler for applications ranging from urban logistics and aerial inspections to environmental monitoring and emergency response. Through cooperative sensing, these UAV networks can significantly enhance angular resolution and achieve high-fidelity environmental imaging, thereby meeting the stringent requirements of safety, reliability, and responsiveness in low-altitude economic activities. For example, the framework in [57] demonstrated how joint optimization can improve target detection in cluttered environments while mitigating multi-user interference, thereby boosting both sensing accuracy and communication reliability. Moreover, ISAC inherently enables implicit channel state information acquisition through environmental sensing, reducing dependence on conventional pilot-based estimation and further improving spectral efficiency [58, 59, 60]. This efficiency gain not only conserves limited spectrum resources but also ensures robust connectivity for diverse services in the LAE.
@@
In addition to its dual communication and sensing capabilities, ISAC fundamentally expands the operational capabilities of UAV networks by enabling advanced spatial awareness and autonomy, which is a key technological driver for the LAE. Applications such as high-precision three-dimensional (3D) mapping, dynamic trajectory planning, and swarm coordination become feasible when sensing and communication are tightly integrated. For instance, [61] introduced a cooperative uplink-downlink reconstruction framework, where base stations performed active sensing in the downlink while passive sensing was realized through base station-user equipment cooperation in the uplink. The fusion of complementary sensing results yielded higher spatio-temporal resolution, broader coverage, and improved reconstruction accuracy. Similarly, environmental data generated via ISAC was shown to enhance real-time collision avoidance and collaborative perception in UAV swarms [62], supporting safer autonomous operations in dense and dynamic LAAs. Advanced fusion strategies, such as the multilevel weighted fusion framework in [63], further improved UAV localization by combining symbol-level and data-level information, thereby enhancing estimation performance. Moreover, [64] proposed a collaborative ISAC-based framework for UAV-terrestrial base station cooperation, achieving high-precision localization alongside reliable communication in scenarios with multiple moving users.
In summary, the integration of ISAC within LAWNs not only addresses the challenges of spectral congestion and payload constraints. It also enables distributed sensing, collaborative perception, and intelligent UAV autonomy. This progression underscores ISAC’s pivotal role in transforming LAWNs into intelligent, resilient, and spectrum-efficient infrastructures that can support mission-critical low-altitude applications such as urban monitoring, disaster response, traffic management, and environmental surveillance.
3.2 Edge intelligence
Edge intelligence integrates AI capabilities into the network edge [65], enabling data processing and inference to be executed directly at edge nodes such as UAVs, base stations, and edge servers. By reducing the reliance on transmitting massive amounts of raw data to remote cloud centers, this paradigm effectively mitigates high latency and alleviates bandwidth congestion, thereby offering a more efficient, scalable, and resilient computing architecture.
In LAWNs, the incorporation of edge intelligence significantly enhances system performance by exploiting the distributed, mobile, and cooperative nature of UAVs. As illustrated in Fig. 4, the evolution of edge intelligence in UAV-enabled networks can be conceptually divided into three progressive stages:
  • Stage 1 (Local computing): In this stage, each UAV executes AI tasks independently using its limited local computation resources. Although suitable for lightweight inference, the restricted computational power and energy capacity often lead to delays in executing complex algorithms, thereby constraining real-time decision-making and autonomous operation.
  • Stage 2 (Edge-enabled collaborative computing): In this stage, UAVs leverage external edge resources by offloading tasks to nearby static nodes such as ground base stations, and terrestrial edge servers. While this approach enables faster response times and more efficient resource utilization compared to isolated computing, the availability of computing resources is still tied to the fixed locations of edge nodes.
  • Stage 3 (LAWN-enabled edge intelligence): In this stage, UAVs evolve into dynamic and flexible computing nodes that collectively form a self-organizing large-scale aerial-ground edge network. Unlike Stage 2, where edge resources are geographically fixed, UAVs in this stage can mobilize and coordinate on demand, migrating toward hotspots such as large-scale events or disaster areas to provide additional communication and computation capacity. AI inference and training are thus performed in a highly adaptive manner through federated learning, proactive task offloading, and mobility-aware resource orchestration. This stage represents a paradigm shift from static edge collaboration to a network-wide, elastic, and resilient intelligence infrastructure empowered by UAV mobility.
Such an evolution is particularly crucial for enabling and sustaining LAE activities, where UAV operations must cope with dynamic and uncertain environments and undertake mission types highly relevant to LAE, such as real-time logistics routing, aerial infrastructure inspection, target tracking, and environmental perception. Traditional cloud-centric AI architectures are often hindered by excessive latency due to long-distance backhaul links, posing challenges for time-critical LAE applications. In contrast, edge intelligence deploys lightweight AI models directly onto UAVs or proximal edge nodes, enabling immediate decision-making and significantly improving responsiveness, autonomy, and mission reliability [66], thereby ensuring efficient, safe, and reliable operations across diverse low-altitude economic scenarios.
Moreover, LAWNs inherently consist of multiple UAVs, making them well-suited for distributed learning paradigms such as federated learning, which can provide scalable, privacy-preserving, and resource-efficient intelligence for LAE operations. By training models collaboratively across UAV nodes without sharing raw data, federated learning not only preserves privacy but also reduces communication overhead [67]. For example, [68] introduced a UAV-enabled mobile edge computing (MEC) framework for air-ground integrated networks, where UAVs acted as aerial base stations to provide adaptive coverage and edge computing support for Internet-of-things (IoT) devices under stringent energy and accuracy constraints. Similarly, [69] proposed a two-tier hierarchical federated learning framework, in which UAVs served as mobile relays that both enhance connectivity to a central server and conducted local computations, thereby improving scalability in large-scale, infrastructure-limited scenarios.
Beyond distributed learning, recent advancements explored broader forms of UAV-enabled edge intelligence that target diverse application domains and performance metrics, many of which are directly applicable to LAE services. In [70], a UAV-assisted IoT data collection framework was introduced under segmented fading channels, achieving lower age of information (AoI), reduced energy consumption, and shorter mission completion time through cooperative UAV operation. In [71], a flying-edge computing framework was proposed for smart agriculture, where UAVs equipped with AI-powered sensors execute tasks such as disease detection and irrigation control directly on-board, enabling real-time decision-making in connectivity-limited areas. Additionally, [72] presented a UAV-assisted MEC framework that jointly optimizes UAV deployment and computation offloading to minimize average response delay, ensuring rapid adaptability to dynamic environments.
In summary, these developments illustrate that edge intelligence, when integrated with LAWNs, evolves into a distributed and cooperative computing infrastructure. It enhances mission responsiveness, privacy preservation, and scalability. By exploiting UAV mobility and the spatial diversity of LAWNs, edge intelligence establishes the foundation for autonomous and resilient aerial networks capable of supporting diverse low-altitude economic applications, including emergency relief, logistics delivery, smart city management, and public safety assurance.

### v2

@@ 2.2 LAWN Channel Modeling
Channel modeling serves as the cornerstone for ensuring safe and reliable UAV operations, as it captures the propagation behaviors that underpin performance evaluation and system design. In this context, spectrum allocations define the operating conditions that directly shape the propagation characteristics. To this end, the International Civil Aviation Organization (ICAO) requires that control and non-payload communication (CNPC) links operate in protected spectrum bands. In line with this requirement, the International Telecommunication Union (ITU) has designated specific portions of the L-band (960–977 MHz) and C-band (5030–5091 MHz) for CNPC services in UAV systems [khuwaja2018survey]. For beyond-line-of-sight (LoS) operations involving satellite communication, the Ku-band and Ka-band have been reserved under the aeronautical safety spectrum framework to support robust satellite-based CNPC connectivity. In addition, national regulators, such as those in China, have authorized supplementary LoS frequency bands to meet the rising demand for UAV deployments in densely utilized low-altitude airspaces [luo]. Based on the allocated bands, traditional studies on aeronautical channel modeling primarily target high-altitude, long-range aircraft operating under transcontinental or military flight conditions. The work in [bello1973aeronautical] introduced one of the earliest statistical frameworks for such environments, analyzing Doppler spread and multipath dispersion caused by rapid airborne motion. In [haas2002aeronautical], the authors categorized modeling approaches into deterministic, stochastic, and geometry-based types, highlighting the influence of aircraft altitude, elevation angle, and antenna orientation on channel characteristics such as path loss and Ricean K-factors. The authors in [jain2012requirements] further examined wireless datalink requirements for large-scale unmanned systems, analyzing the trade-offs among various candidate technologies under high-mobility and limited-spectrum conditions. While primarily focused on conventional high-altitude aircraft, the resulting models and statistical parameters offer a foundational reference for extending channel characterization into LAWNs.
@@
The A2G channel is pivotal for facilitating data exchange within LAWNs, making accurate modeling of the A2G propagation channel essential for the design of resilient drone-based communication systems. In urban environments where LAWNs are commonly deployed, the propagation channels are often shaped by dense building structures, irregular street canyons, and pronounced multi-path effects, leading to significant challenges for traditional propagation models. A representative statistical modeling approach is presented in [6492100], where the authors conducted extensive UAV field measurements at low elevation angles across urban environments. They developed a time-series generator capable of capturing the elevation-dependent fading characteristics, providing a realistic stochastic representation of mobile A2G links for simulation and system design. Furthermore, the authors in [xu20223] studied 6G low-altitude mmWave massive multiple-input multiple-output (MIMO) A2G channels in urban scenarios, capturing space-time-frequency non-stationarity with UAV-specific parameters. To further quantify large-scale shadowing behavior of A2G channels, an elevation-dependent shadowing model was proposed in [4483593]. Originally developed for high-altitude platforms (HAPs), the model remains applicable to LAWNs operating over urban areas. The key insight is that the shadowing loss exhibits a strong dependence on the elevation angle, following an exponential trend as the UAV moves closer to the horizon, thus affecting LoS probability and average signal strength. The work of [7037248] investigated an analytical A2G path loss model that integrates stochastic geometry with ray-tracing results, which separates LoS and non-LoS (NLoS) components and incorporates a probabilistic LoS function based on elevation angle and building distribution. Beyond classical stochastic and geometry-based models, data-driven channel prediction offers a complementary way to address the strong nonstationarity of low-altitude links. In the hybrid framework of [8663422], a physics-based state evolution model captures large-scale effects such as elevation-dependent path loss and shadowing, while a learned predictor trained on measurement data models small-scale fading, blockage events, and environment-specific dynamics that are difficult to describe analytically. The combined use of model-based state evolution and data-driven learning is particularly suitable for LAWNs, enabling accurate short-term channel forecasting for UAV relays operating in complex urban environments.
@@
In addition to urban deployments, the study in [8048502] proposed an analytical path-loss model for suburban A2G transmissions. By incorporating UAV height and LoS probability, it facilitates efficient planning of UAV trajectories and resource allocation in semi-urban environments. Regarding over-water scenarios, the authors in [5089553] developed a geometry-based multi-path model capturing sea surface reflections and long-range propagation delays for A2G links. Meanwhile, the work in [5741881] conducted C-band measurements over sea surfaces at low altitudes, characterizing channel fading patterns and multi-path effects essential for UAV telemetry and maritime missions. Notably, LAWN systems are increasingly integrated into cellular networks, which exhibit distinct propagation characteristics. The work in [7936620] conducted empirical studies of UAV communication over commercial long-term evolution (LTE) networks, which revealed that drones flying above the downtilted beam of base station antennas are likely to result in higher path loss and interference levels. The authors in [8606910] further explored mid-band 5G service delivery for ultra-low-altitude UAV base stations, whose measurements demonstrated significant Doppler effects and elevation-dependent delay profiles, emphasizing the importance of precise modeling for 5G aerial deployments. To complement theoretical modeling, a series of wideband channel measurement campaigns were conducted in [matolak2017air1, sun2017air2, matolak2017air3], targeting A2G links over diverse terrains including over-water, mountainous, and suburban areas, providing empirical models capturing delay spread, angular dispersion, and Ricean fading statistics with varying antenna heights and aircraft velocities.
Table 4: Summary of representative LAWN channel modeling
Scenario Frequency band Link Aircraft type Ref.
Urban (100–170 m altitude) 2 GHz A2G Airship [6492100]
Dense urban MmWave A2G/A2A Low-altitude UAVs [xu20223]
Urban Sub-6 GHz A2G HAP [4483593, 7037248]
Over-water / maritime C band A2G Small UAVs [5089553, 5741881]
Rural/mountain/near-urban L/C bands A2G UAVs [matolak2017air1, sun2017air2, matolak2017air3]
Suburban / rural LTE band A2G UAVs [7936620]
Open field 2.4/5 GHz (IEEE 802.11) A2G Fixed-wing UAVs [4086461]
Terrestrial and maritime C band A2A Small fixed-wing UAVs [7991486]
Aerial relay Sub-6 GHz A2A UAVs [dac.1212]
Open field/sea VHF/L/C bands Air-to-space/A2A/A2G Civil aircraft [bello1973aeronautical, haas2002aeronautical, jain2012requirements]
*VHF: very high frequency (30–300 MHz); L-band (960–977 MHz); C-band (5030–5091 MHz)
@@
3 LAWN Performance Metrics
In this section, we investigate key performance metrics for LAWN designs, focusing on latency, energy efficiency (EE), and scalability.
3.1 Timeliness and Latency
In practice, LAWNs are frequently deployed to support time-sensitive applications, where the transmission delay may lead to severe operational and safety implications. To name a few, in emergency response, the timely delivery of situational data such as survivor locations, fire spread, or hazardous material detection is essential for enabling rapid decision-making and coordinated rescue actions. Therefore, one of the core design principles for LAWNs is the minimization of end-to-end latency while maintaining reliable data transmission. To achieve this, finite-block length (FBL) transmission has emerged as an important technique for managing communication in constrained networks. Unlike the traditional Shannon capacity framework, which assumes infinitely long codewords, FBL transmission provides a more practical characterization of the achievable rates given block lengths and error probability. More importantly, the channel conditions in LAWNs often change rapidly, leading to the reliable infinite block length transmission assumption being invalid. The authors of [zhou2023finite, zhou2017second, zhou2019non] investigated finite blocklength multiterminal lossy source coding, providing new benchmarks for efficient compression with multiple encoders, decoders, and source sequences, thereby strengthening the theoretical foundation of multiterminal low-latency communications and greatly promoting the development of low-altitude low-latency networking. The work of [9126811] investigated IoT service provisioning by leveraging FBL transmission in reducing communication delays within drone-aided networks, with drones being both communication relays and computational resources. Moreover, Jin et al. proposed the co-design of sensing, communication, and control in LAWNs, where FBL transmission is utilized with remote control tasks to minimize latency. The results show that optimizing resource allocation for communication and control can lead to lower delays and better overall performance [11045436].
Beyond FBL transmission, some recent studies focus on low-latency communications. For instance, the authors in [liu2023predictive] presented a predictive precoder design for ultra-reliable low-latency communication (URLLC) systems, where a deep learning (DL)-based framework was developed to predict the precoder via historical channel data. By leveraging a convolutional long short-term memory (CLSTM) network, reliable communication with minimal latency can be guaranteed, which demonstrates superior frame error rate (FER) performance approaching the theoretical lower bound with reduced latency. More relatively, the work of [cai2022resource] explored resource allocation for a two-way UAV relaying system under URLLC constraints, focusing on minimizing latency in both the forward and backward links. An iterative approach was proposed to maximize the system throughput while fulfilling the latency and decoding error probability requirements.
Moreover, the age of information (AoI) is another critical metric facilitating delay-aware LAWN design. In particular, the AoI measures the freshness of data in a communication network, where outdated data can lead to poor decision-making performance. Toward this end, the study in [9151993] explored UAV-assisted wireless-powered IoT networks, which optimized UAV trajectories to minimize AoI and ensure that the sensor data remains fresh. Furthermore, Gao et al. examined AoI-sensitive data collection in multi-UAV-assisted wireless sensor networks [10007855], in which a clustering algorithm, devised to optimize both AoI and energy consumption in multi-UAV systems, was developed. The results show that by adjusting the UAV flight paths and clustering strategies, AoI can be minimized such that LAWN nodes have access to the freshest data with low energy usage, highlighting that AoI management is essential for real-time collaboration and successful task execution within LAWNs. The importance of real-time scheduling in LAWN systems was studied in [10978356] to manage communication effectively. By prioritizing the transmission of high-priority data and optimizing the transmission schedules, real-time scheduling can help minimize latency and AoI.
3.2 Energy-Efficient Design
The sustainability of LAWNs is inherently constrained by the limited onboard energy and computational resources of UAVs, making energy-efficient design essential for long-term viability in real-world applications. Early studies primarily examined energy-aware deployment strategies to extend system lifetime. For instance, [9126212] proposed energy-saving deployment algorithms for UAV swarms to optimize wireless coverage while minimizing energy consumption. In a similar approach, the study of [8364586] addressed the rapid deployment problem of heterogeneous UAVs to provide wireless coverage over a target area, where the objective for minimizing the maximum deployment delay to ensure fairness and minimizing the total deployment delay to enhance efficiency were addressed, respectively. However, these works largely focus on stationary optimization and do not fully exploit the high-mobility characteristics of LAWNs. In the context, the authors in [8068199] investigated joint trajectory and transmit power control for UAV relay networks to minimize the outage probability. By carefully devising UAV trajectories and power allocation, energy consumption can be reduced without sacrificing communication reliability. Further, the study in [yang2019energy] proposed a UAV-assisted backscatter communication system for energy-efficient IoT, which shows a tradeoff between collection location and outage probability on the improvement of EE. On this basis, the work of [11129138] investigated the trade-off between communication/sensing performance and energy consumption in ISAC-LAWN systems, where the goal was to maximize worst-case EE by jointly designing beamforming and UAV trajectory.
Note that [9126212, 8364586, 8068199, yang2019energy, 11129138] primarily focused on the transmit power control, significantly lower than the propulsion-related power. Hence, the work of [10159441] proposed a LAWN system simultaneously sensing ground users and forwarding the sensed information. The authors formulated a multi-objective optimization problem to maximize the EE by jointly optimizing user scheduling, movement-related power, and UAV trajectory. While existing methods have improved network EE, their optimization objectives are often limited to reducing the energy consumption of individual UAVs, overlooking the potential gains from cooperative flight dynamics in LAWNs. In real-world deployments, UAVs inevitably interact with surrounding airflow patterns, which can be strategically exploited to boost overall EE. Specifically, airflow over UAV wings generates wake vortices that create an upwash region behind the wingtips. A trailing UAV entering this region experiences reduced aerodynamic drag, thereby lowering propulsion energy consumption, as well-documented in avian flocking. Building on this principle, the studies in [weimerskirch2001energy, wu2025toward] introduced an ISAC-LAWN design that jointly takes into account aerodynamic factors and sensing/communication performance, showing that a ’V’-shaped formation is likely to be aerodynamically favorable.
It is evident that the above approaches can only save energy to extend network lifetime, lacking the ability to recharge the battery to achieve a long-term power supply. To address this challenge, the authors in [10139785] explored the performance improvement of UAV communication systems integrated with solar energy harvesting. The uncertainty of solar energy supply was investigated, followed by presenting an iterative optimization approach to maximize throughput. The simulations show that integrating solar energy into LAWN systems can significantly extend the operational duration, thereby enhancing the sustainability in remote areas where recharging infrastructure is scarce. Meanwhile, the work of [8941314] investigated energy-efficient LAWN communications design with energy harvesting, where trajectory optimization techniques were used to minimize energy consumption while maximizing the coverage area as well as maintaining QoS.
3.3 Scalability and Connectivity
Benefiting from the high mobility and on-demand deployment capabilities, scalability and connectivity are essential for LAWNs to adapt to various practical requirements. The performance of large-scale LAWN networks is heavily influenced by the ability to manage communication capacity, coverage area, and resilience under dynamic conditions, such as UAV movement and interference. For instance, to reveal the fundamental performance limits of beam alignment for mobile targets such as UAVs in low-altitude networks, the authors in [zhou2021resolution, zhou2022resolution] investigated mmWave multiple-antenna communications, deriving theoretical benchmarks for optimal non-adaptive query procedures in multidimensional target search and extending them to scenarios with multiple and moving targets. They further demonstrated the benefits of adaptive query procedures, thereby providing finite-blocklength benchmarks that enable reliable connectivity in low-altitude networks. As evidenced by [10557642], wherein network localization and formation control in LAWN systems with asynchronous agents were investigated. Specifically, the authors developed a message-passing algorithm using Gaussian messages for efficient localization and clock offset estimation to maintain robust information exchange across all LAWN nodes. Subsequently, two types of control policies, namely angle-based and distance-based formation control, were employed for adaptive obstacle avoidance. By doing so, the LAWN systems are capable of quickly responding to environmental changes, and thereby, enhancing scalability and connectivity.
However, the coexistence of heterogeneous LAWN nodes sharing spectral and spatial domains creates a complex interference environment that challenges large-scale deployment and reliable interconnection. The main challenges stem from beam pattern overlaps and waveform interactions across both spatial and temporal dimensions. To tackle these issues, the work in [WU2025103451] introduced a mode-switching strategy, where UAVs alternate between jamming and information transmission. By optimizing mode selection, scheduling, and trajectories, this approach ensures strong A2G links even in the presence of mobile ground users. Furthermore, [9143143] proposed a deep learning–based predictive beamforming scheme that mitigated UAV jittering by forecasting the angles between UAV and user, thereby enabling robust and reliable UAV communications. Moreover, the authors in [9551699] proposed a dynamic interference management framework by jointly optimizing UAV trajectories and power allocation, thereby improving overall network efficiency under dense deployments. Beyond intra-LAWN coordination, such solutions can also be extended to satellite integration. For example, the authors in [10680056] designed robust beamforming schemes for SAGIN to guarantee reliable downlink communications.
@@ 4.1 Integrated Sensing and Communications
ISAC represents a transformative paradigm that unifies wireless communications and environmental sensing into a single framework [liu2022integrated, mao1, mao2, wu2025toward1, mao3, mao4]. By adopting a co-design architecture, ISAC eliminates the need for deploying separate sensing or communication subsystems on aerial platforms within LAWNs. This integration not only reduces the spectral footprint, hardware redundancy, and energy consumption, but also enhances the agility of the system. For instance, a UAV transmitting communication signals to ground users can simultaneously leverage the backscattered echoes for obstacle detection or trajectory tracking of neighboring UAVs [jiang2024uav], thereby avoiding reliance on dedicated radar or LiDAR sensors.
@@
  • Stage 2 (ISAC systems): In this stage, each UAV integrates sensing and communication functionalities into a unified platform by sharing hardware and spectral resources. This design improves spectral efficiency, reduces payload burden, and enhances cost-effectiveness compared to separated systems. However, its benefits remain limited to individual UAVs, and scalability across large networks is not yet fully achieved.
@@
This evolutionary trajectory highlights the maturation of ISAC from a dual-functional system into a network-wide intelligent infrastructure, which is particularly crucial for supporting the burgeoning LAE. In such scenarios, UAV networks serve as a vital enabler for applications ranging from urban logistics and aerial inspections to environmental monitoring and emergency response. Through cooperative sensing, these UAV networks can significantly enhance angular resolution and achieve high-fidelity environmental imaging, thereby meeting the stringent requirements of safety, reliability, and responsiveness in low-altitude economic activities. For example, the framework in [wang2025transmit] demonstrated how joint optimization can improve target detection in cluttered environments while mitigating multi-user interference, thereby boosting both sensing accuracy and communication reliability. Moreover, ISAC inherently enables implicit channel state information acquisition through environmental sensing, reducing dependence on conventional pilot-based estimation and further improving spectral efficiency [pyo2025robust, 10791452, 10543024]. This efficiency gain not only conserves limited spectrum resources but also ensures robust connectivity for diverse services in the LAE.
@@
In addition to its dual communication and sensing capabilities, ISAC fundamentally expands the operational capabilities of UAV networks by enabling advanced spatial awareness and autonomy, which is a key technological driver for the LAE. Applications such as high-precision three-dimensional (3D) mapping, dynamic trajectory planning, and swarm coordination become feasible when sensing and communication are tightly integrated. For instance, [lu2024deep] introduced a cooperative uplink-downlink reconstruction framework, where base stations performed active sensing in the downlink while passive sensing was realized through base station-user equipment cooperation in the uplink. The fusion of complementary sensing results yielded higher spatio-temporal resolution, broader coverage, and improved reconstruction accuracy. Similarly, environmental data generated via ISAC was shown to enhance real-time collision avoidance and collaborative perception in UAV swarms [wang2024cooperative], supporting safer autonomous operations in dense and dynamic LAAs. Advanced fusion strategies, such as the multilevel weighted fusion framework in [zhang2025novel], further improved UAV localization by combining symbol-level and data-level information, thereby enhancing estimation performance. Moreover, [hu2025collaborative] proposed a collaborative ISAC-based framework for UAV-terrestrial base station cooperation, achieving high-precision localization alongside reliable communication in scenarios with multiple moving users. Plugging ISAC into LAWNs not only addresses the challenges of spectral congestion and payload constraints. It also enables distributed sensing, collaborative perception, and intelligent UAV autonomy. This progression underscores ISAC’s pivotal role in transforming LAWNs into intelligent, resilient, and spectrum-efficient infrastructures that can support mission-critical low-altitude applications such as urban monitoring, disaster response, traffic management, and environmental surveillance.